OpenStudy (anonymous):

A farmer wants to fence an area of 6 million square feet in a rectangular field and then divide it in half with a fence parallel to one of the sides of the rectangle. How can he do this so as to minimize the cost of the fence?

OpenStudy (anonymous):

you have to use 2 equations. Since its a rectangle and you know the area u can say the length is y and the width is x. Therefore the equation for the area is xy=6(10^6). Next you know that the amount of fence you are gonna use is equal to the perimeter and the fence that splits it into 2. so P=2(x+y) + x from the first equation solve for y: y=6(10^6)/x and input into 2nd equation: P=3x+12(10^6)/x Differentiate: P'=3-12(10^6)/(x^2) After a little rearrangement P'=(3x^2-12(10^6))/(x^2) which yields the points x=-2000,0,2000 reject the negative and 0 value leaving x to b 2000. Solve for y in the first equation and you'll get that y =3000